OpenStudy (anonymous):

a length of a rectangle is 5 inches more than the width. the area of the rectangle is 84 sq. inches. find the width of the rectangle a length of a rectangle is 5 inches more than the width. the area of the rectangle is 84 sq. inches. find the width of the rectangle @Mathematics

OpenStudy (anonymous):

L=W+5 LW=84 W=84/L W=84/(W+5) W^2+5W-84=0 Quadratic Equation (find the positive root) W=(-5+19)/2 W=7

OpenStudy (anonymous):

thank you!